OpenStudy (anonymous):

the half-life of radium 226 is 1590 years. a) if a sample has a mass of 150 mg find the mass that will remain after 1000 years. A=150(1/2)^(1000/1590)= 96.9 mg remain. b) after how many years will only 50 mg remain???

OpenStudy (campbell_st):

well you need to find t so the equation sets up as \[50 = 150\times (\frac{1}{2})^{\frac{t}{1590}}\] just solve for t you'll need logs...

OpenStudy (anonymous):

when do the logs come in? i tried the problem but did it wrong